Driving down a mountain, Tom finds that he has descended 1800 ft in elevation by the time he is 3.25 miles horizontally away from the top of the mountain. Find the slope of his descent to the nearest hundredth.
Answer by stanbon(57228)

Driving down a mountain, Tom finds that he has descended 1800 ft in elevation by the time he is 3.25 miles horizontally away from the top of the mountain. Find the slope of his descent to the nearest hundredth.
---------------
vertical change = 1800 ft
horizontal change = 3.25*5280 ft = 17160 ft.
--------------------
slope = vert/hor = 1800/17160 = 0.1049
=================
Cheers,
Stan H.